# Impact of space-time curvature coupling on the vacuum energy induced by a magnetic topological defect in flat space-time of arbitrary dimension

### v1

@@ 5 Summary
In flat $(d+1)$-dimensional space-time, we consider the vacuum polarization of a quantized charged scalar field in the presence of a magnetic topological defect, modeled as an impenetrable tube of finite thickness carrying magnetic flux. The induced vacuum energy is studied with particular attention to its dependence on the coupling parameter $(\xi)$, which characterizes the interaction of the scalar field with space-time curvature. The scalar matter field is assumed to be excluded from the tube interior and to satisfy generalized Robin-type boundary conditions (5) on its surface. These boundary conditions are specified by a single parameter, $\pi/2\leq\theta<\pi/2$. The special cases $\theta=0$ and $\theta=-\pi/2$ correspond to Dirichlet and Neumann boundary conditions, respectively.
@@
Our analysis demonstrates that, in flat space-time, the induced vacuum energy within the plane transverse to the tube $E^{(d+1)}=E^{(d+1)}_{can}+(1/4-\xi)E^{(d+1)}_{\xi}$ is independent of the coupling $\xi$ of the scalar field’s interaction with the space-time curvature only in the special cases of Dirichlet and Neumann boundary conditions. But for general Robin boundary conditions, the induced vacuum energy in flat space-time depends on the coupling $\xi$ and exhibits a nontrivial dependence on the parameter of the boundary condition $\theta$.
@@
We computed $E^{(d+1)}_{\xi}$, which determines the dependence of the total induced vacuum energy on the coupling $\xi$ for the cases of $d=2,3,4$ and negative values of boundary parameter $\pi/2\leq\theta\leq 0$, see Figs.1 and 2. As one can see, the vacuum effects increase strongly as the tube radius decreases. It is also worth noting that as the thickness of the tube decreases, the value of the boundary condition parameter at which the function reaches its maximum shifts toward the Neumann boundary condition, $\theta=-\pi/2$. Importantly, the induced energy increases with the space-time dimensionality and for sufficiently small transverse sizes of the topological defect $mr_{0}$.
@@
We reconstruct the dependence of the induced dimensionless vacuum energy $E^{(d+1)}_{can}/m^{d-1}$ on the tube thickness for the Dirichlet and Neumann cases, see Fig.3, and compare it with the maximal value of $E^{(d+1)}_{\xi}/m^{d-1}$ attained at some $\theta_{\mathrm{max}}$, see Table 1. We conclude that for certain values of the boundary condition parameter $\pi/2<\theta<0$, the quantity $E^{(d+1)}_{\xi}/m^{d-1}$ is of the same order of magnitude, or even larger, than $E^{(d+1)}_{can}/m^{d-1}$ for the same $\theta$. This means that the $E^{(d+1)}_{\xi}$ plays an important role in determining the induced energy, and its dependence on the parameter $\xi$ is significant. We conclude that careful measurements of vacuum polarization effects in flat space-time may, in principle, provide an independent way to probe the $\xi$ coupling that determines the interaction of a scalar field with space-time curvature.
@@
When the boundary condition parameter $\theta$ takes positive values, the problem becomes particularly interesting and requires further thorough analysis. In this case, bound-state solutions of the Fock–Klein–Gordon equation may contribute to vacuum polarization, which can result in nontrivial discontinuous dependence of vacuum effects on $\theta$, similar to the behavior observed for the induced magnetic flux [42].

### v2

@@ 4 Induced vacuum energy in higher-dimensional spaces
5 Asymptotic behavior of the induced $\xi$-dependent energy for thin and thick tube limits
In this section, we will closely examine the dependence of the induced vacuum energy $E_{\xi}^{(d+1)}$ on the thickness of the tube, namely on the parameter $mr_{0}$.
The study of the asymptotic behavior of the $\xi$-dependent induced vacuum energy in the limits of very thin and very thick magnetic tubes is motivated by the need to clarify how the coupling to space-time curvature manifests itself in extreme physical regimes. In the thin-tube limit, the magnetic tube effectively reduces to a zero-thickness string, while the thick-tube limit probes the suppression of vacuum effects under the increasing of the tube radius $r_{0}$. On the other hand, the thickness of a topological defect can be estimated from the approximate condition $m_{H}r_{0}\sim 1$, where $m_{H}$ is the scale of spontaneous symmetry breaking (or the mass of the corresponding Higgs field) responsible for the formation of the topological defect in astrophysical or condensed-matter systems. This relation implies $mr_{0}\sim m/m_{H}$. Therefore, the consideration of thin and thick tube limits effectively corresponds to studying vacuum polarization effects for light and heavy scalar fields relative to the symmetry-breaking scale.
For $mr_{0}\ll 1$, we get from (30)
$\frac{E^{(d+1)}_{\xi}}{m^{d-1}}=\frac{C(F,d,\theta)}{x_{0}^{d-1}},\qquad C(F,d,\theta)=2\frac{(4\pi)^{1-d/2}}{\Gamma\left(\frac{d-2}{2}\right)}\int\limits_{0}^{\infty}dv\,v^{d-2}\,\mathcal{D}_{\xi}(F,\theta,v).$ (34)
Unfortunately, the coefficient $C(F,d,\theta)$ cannot be written in analytical form. But numerical computations show that it is a decreasing function of $d$, which can be roughly approximated for $d\leq 8$ as
$C=e^{-a-b\,d},$ (35)
where $a$, $b$ are positive coefficients of $F$ and $\theta$ parameters. Then at $mr_{0}\ll 1$ we get
$\frac{E^{(d+1)}_{\xi}}{m^{d-1}}=e^{\ln x_{0}-a}e^{-d(b+\ln x_{0})}=\frac{\beta}{(\gamma x_{0})^{d-1}},\quad mr_{0}\ll 1,$ (36)
where $\beta=e^{-(a+b)}$, $\gamma=e^{b}$. Such a simple representation makes it possible, at a qualitative level, to understand the dependence (obtained numerically) of the induced vacuum energy on the tube thickness and on the space dimension. By analyzing the function in the exponent, one can see that for a magnetic tube of small thickness [^1] If we put it more precisely, the function $C(d)$ has the form $C=e^{\alpha+\beta\,d+\gamma/d+\delta\gamma/d^{2}}$ and $mr_{c}=e^{\beta-{2\delta}/{d^{3}}-{\gamma}/{d^{2}}}$. For example, for $F=1/2$ and $\theta=-1$, the coefficients are $\alpha=-8$, $\beta=-1.78$, $\gamma=20.48$, and $\delta=-14.2$, which, in turn, give $mr_{c}=0.05$ for $d=3$ and $mr_{c}=0.073$ for $d=4$., $x_{0}<mr_{c}$, $mr_{c}=e^{-b}$, the induced vacuum energy increases as the space dimension $d$ grows, whereas for a sufficiently large tube thickness, the induced vacuum energy instead decreases with increasing $d$. For example, for $F=1/2$ and $\theta=-1$, the coefficients are $a=0.55$ and $b=2.48$, which, in turn, give $mr_{c}=0.084$ and $\beta=0.048$, $\gamma=11.94$.
For $mr_{0}\gg 1$, we get from (30)
$E^{(d+1)}_{\xi}=\frac{2}{r_{0}^{d-1}}\frac{(4\pi)^{1-d/2}}{\Gamma\left(\frac{d-2}{2}\right)}\int\limits_{x_{0}}^{\infty}dv\,v^{2}\left(v^{2}-x_{0}^{2}\right)^{\frac{d-4}{2}}\,\mathcal{D}^{as}_{\xi}(F,\theta,v),$ (37)
where $\mathcal{D}^{as}_{\xi}$ is the asymptotic form of the function $\mathcal{D}_{\xi}$ (32) at large $mr_{0}$. Numerical computations show that it can be approximated as
$\mathcal{D}^{as}_{\xi}=\frac{e^{-\bar{a}-\bar{b}\,x_{0}}}{x_{0}},\quad x_{0}\gg 1,$ (38)
where $\bar{a}$, $\bar{b}$ are positive coefficients of $F$ and $\theta$ parameters. By substituting the obtained asymptotic expression into (37), one can carry out the integration explicitly and obtain
$\frac{E^{(d+1)}_{\xi}}{m^{d-1}}=\frac{2\bar{b}}{e^{\bar{a}}(2\pi\bar{b}\,x_{0})^{\frac{d-1}{2}}}K_{\frac{d-1}{2}}(\bar{b}\,x_{0}){\approx\frac{\bar{b}\,e^{-\bar{a}}}{(2\pi)^{\frac{d}{2}-1}}\frac{e^{-\bar{b}\,x_{0}}}{(\bar{b}x_{0})^{\frac{d}{2}}},}\quad mr_{0}\gg 1,$ (39)
where $K_{\mu}(u)$ is the Macdonald function. For example, for $F=1/2$ and $\theta=-1$, the coefficients are $\bar{a}=5.22$ and $\bar{b}=7.23$.
To conclude this section, it is important to note that variations in the tube radius in physical systems may be accompanied by changes in the surface properties of the tube and the corresponding boundary parameter $\theta$, see, e.g., [28, 8].
@@ 6 Summary
In this paper, we considered the possibility of the manifestation of coupling $\xi$ of the scalar field’s interaction with the space-time curvature in careful measurements of vacuum polarization effects in flat space-time.
A nonzero value of the coupling $\xi$ of the scalar field’s interaction with the space-time curvature modifies the effective mass of particles, which affects particle creation and decay processes, influences the inflationary stage of the Universe’s evolution, the stability of the electroweak vacuum, and the spectrum of Hawking radiation from black holes for scalar fields. At the same time which exact value of the coupling constant is realized in nature remains an open question. An experimental determination is hindered by the typically negligible value of the Ricci scalar $R$ for most cosmological and astrophysical problems. In fact, the contribution of the curvature coupling constant becomes significant when the scalar field is in regions with large curvature, such as in the early Universe (during inflationary or radiation-dominated epochs) or near strong gravitational sources (black holes, neutron stars). Therefore, studies of the possible manifestation of coupling $\xi$ in flat space-time are of considerable interest.
In flat $(d+1)$-dimensional space-time, we consider the vacuum polarization of a quantized charged scalar field in the presence of a magnetic topological defect, modeled as an impenetrable tube of finite thickness carrying magnetic flux. The induced vacuum energy is studied with particular attention to its dependence on the coupling parameter $(\xi)$, which characterizes the interaction of the scalar field with space-time curvature. The scalar matter field is assumed to be excluded from the tube interior and to satisfy general Robin-type boundary conditions (5) on its surface. These boundary conditions are specified by a single parameter, $\pi/2\leq\theta<\pi/2$. The special cases $\theta=0$ and $\theta=-\pi/2$ correspond to Dirichlet and Neumann boundary conditions, respectively.
@@
Our analysis demonstrates that, in flat space-time, the induced vacuum energy within the plane transverse to the tube $E^{(d+1)}=E^{(d+1)}_{can}+(1/4-\xi)E^{(d+1)}_{\xi}$ is independent of the coupling $\xi$ of the scalar field’s interaction with the space-time curvature only in the special cases of Dirichlet and Neumann boundary conditions $(E^{(d+1)}_{\xi}=0)$. But for general Robin boundary conditions, the induced vacuum energy in flat space-time depends on the coupling $\xi$ $(E^{(d+1)}_{\xi}\neq 0)$ and exhibits a nontrivial dependence on the parameter of the boundary condition $\theta$.
@@
We computed $E^{(d+1)}_{\xi}$, which determines the dependence of the total induced vacuum energy on the coupling $\xi$ for the cases of $d=2,3,4$ and negative values of boundary parameter $\pi/2\leq\theta\leq 0$, see Figs.1 and 2. As one can see, the vacuum effects increase strongly as the tube radius decreases. It is also worth noting that as the tube thickness decreases, the value of the boundary condition parameter at which the function reaches its maximum shifts toward the Neumann boundary condition, $\theta=-\pi/2$. Importantly, the induced energy increases with increasing the space-time dimensionality and for sufficiently small transverse sizes of the topological defect $mr_{0}<mr_{c}$, but decreases with increasing dimensionality at large $mr_{0}>mr_{c}$, where $mr_{c}\approx 0.05$ for $d=2,3,4$ and grows slowly as the spatial dimension increases.
@@
We reconstruct the dependence of the induced dimensionless vacuum energy $E^{(d+1)}_{can}/m^{d-1}$ on the tube thickness for the Dirichlet and Neumann cases, see Fig.3, and compare it with the maximal value of $E^{(d+1)}_{\xi}/m^{d-1}$ attained at some $\theta_{\mathrm{max}}$, see Table 1. We conclude that for certain values of the boundary condition parameter $\pi/2<\theta<0$, the quantity $E^{(d+1)}_{\xi}/m^{d-1}$ is of the same order of magnitude, or even larger, than $E^{(d+1)}_{can}/m^{d-1}$ for the same $\theta$. This means that the $E^{(d+1)}_{\xi}$ plays an important role in determining the induced energy, and the dependence of the induced vacuum energy (33) on the parameter $\xi$ is significant.
We conclude that careful measurements of vacuum polarization effects in flat space-time may, in principle, provide an independent way to probe the $\xi$ coupling that determines the interaction of a scalar field with space-time curvature. Since the induced vacuum energy effect is strongest for tubes with $mr_{0}\ll 1$, efforts to observe the $E_{\xi}$ contribution should focus on satisfying this requirement. Additionally, it is essential to establish conditions corresponding to the value of the boundary parameter $\theta$ that maximizes $E_{\xi}$.
It should be noted that the results obtained here for the physically interesting $(3+1)$-dimensional space–time, as well as for higher-dimensional spaces, were derived by performing the corresponding integration of the induced vacuum energy expression in $(2+1)$-dimensional space–time, see (33). The results for the induced vacuum energy generated by a magnetic tube with a general Robin boundary condition in $(2+1)$-dimensional space–time were first obtained in [20]. In that paper, the induced vacuum energy was studied only for a fixed thickness of the magnetic tube (ring), $mr_{0}=1/100$, and for various negative values of the Robin boundary parameter $\theta$. In the present paper, we compute the induced energy for an arbitrary thickness of the magnetic tube in $(2+1)$-dimensional space–time, which makes it possible to carry out the integration procedure used in [20] and thus extend the previous results to space–times of arbitrary dimensionality $d$ and arbitrary values of the transverse size of the magnetic tube $mr_{0}$.
@@
When the boundary condition parameter $\theta$ takes positive values, the problem becomes particularly interesting and requires further thorough analysis. In this case, bound-state solutions of the Fock–Klein–Gordon equation may contribute to vacuum polarization, which can result in nontrivial discontinuous dependence of vacuum effects on the tube thickness and $\theta$, similar to the behavior observed for the induced magnetic flux [42].